OpenStudy (anonymous):

Find the coeffcient a of the given term in the expansion of the binomial binomial= (x-2y)^10 term= ax^8y^2

OpenStudy (rational):

\[\large \binom{10}{2} x^{8}(-2y)^2\]

OpenStudy (anonymous):

ok uh how do you do this?

OpenStudy (rational):

http://www.wolframalpha.com/input/?i= \binom{10}{2}+x^{8}%28-2y%29^2

OpenStudy (anonymous):

on the test we are not allowed to use a calculator

OpenStudy (rational):

heard of "binomial coefficient" before ?

OpenStudy (anonymous):

uh sort of but idk how to use it

OpenStudy (anonymous):

wait ok what's the formula that you used?

OpenStudy (rational):

\[\large \binom{n}{k} = \dfrac{n!}{k!(n-k)!}\]

OpenStudy (rational):

plugin \(n=10\) and \(k=2\)

OpenStudy (anonymous):

but the answer is 45 when you do that

OpenStudy (anonymous):

the real answer is suppose to be 180 though

OpenStudy (rational):

Yep! the term is \[\large \binom{10}{2} x^{8}(-2y)^2\] plugin the value of \(\binom{10}{2}\) and simplify

OpenStudy (anonymous):

wait what???? plus in 45 to what???

OpenStudy (rational):

\[\large \binom{10}{2} x^{8}(-2y)^2 = 45 x^8(-2y)^2 = ?\]

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhh wait, I think I get it

OpenStudy (anonymous):

it's 180

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

took me a while

OpenStudy (anonymous):

hey imma write a testimony btw, is it ok if it's like not the best?

OpenStudy (anonymous):

like if it's a couple of words? that ok? or you don't want me to write a sloppy testimony?

OpenStudy (rational):

thats sweet <3

OpenStudy (anonymous):

i guess i'll take that as a yes

OpenStudy (anonymous):

thanks btw

OpenStudy (rational):

yw :) ty for the awesome testimonial XD

OpenStudy (sepeario):

Very well explained.